OpenStudy (anonymous):

solve for M2 using (M1)(V1)= (M2)(V2) (40ppm)(.005L) = M2(.050L) answer must be in ppm

OpenStudy (jfraser):

solve the equation for M2 algebraically. The volumes will cancel, so the answer will automatically have units of ppm.\[M_1*V_1 = M_2*V_2\] so \[M_2 = \frac{M_1 * V_1}{V_2}\] and \[M_2 = \frac{40ppm*0.005L}{0.050L}\] see how the units of volume cancel, and all you're left with is a concentration, in ppm.